OpenStudy (rjtrain1):

2.680+34.200+20.386 @Will.H

OpenStudy (rjtrain1):

\[will- medal\]

OpenStudy (rjtrain1):

sorry but that's wrong

OpenStudy (rjtrain1):

will medal if chance its right

OpenStudy (rjtrain1):

there's 3 numbers to add to each other

Awolflover1 (awolflover1):

we'll get there

Awolflover1 (awolflover1):

once you add that ^ we'll add 20.386 lol

OpenStudy (rjtrain1):

I would whether the answer please

Awolflover1 (awolflover1):

sorry i can't give direct answers.

OpenStudy (rjtrain1):

ok and how do you menchen someone in a quistion

Awolflover1 (awolflover1):

@awolflover1 like that with a ''@''

OpenStudy (rjtrain1):

thank you

Awolflover1 (awolflover1):

you're welcome

OpenStudy (will.h):

2.680+34.200+20.386 = ? forget about the decimal for awhile let's add the constants 2 + 34 + 20 = 56 now add the tenth 6 + 2 + 3 = 11 (can't leave it that way you will have to add 1 to the constants) so then tenth now would be 1 and the constants would change from 56 to 57 now the hundredths: 8 + 0 + 8 = 16 again can't leave it that way we will have to add 1 to the tenths and leave it 6) so hundredths would be 6 and tenths would be now 2 thousandths: 0 + 0 + 6 = 6 So the final answer would be 57.266 hope it helps